Solve the following system of ODE

x'+x-y = e^2t
1
Expert's answer
2020-10-12T13:22:47-0400

"\textsf{Case 1.}\\

\textsf{If} \, y(t) = x(t) \\

x' + x - y = e^{2t}\\

\displaystyle\frac{\mathrm{d}x}{\mathrm{d}t} + x - y = e^{2t}\\

\textsf{If}\hspace{0.1cm} y = x = e^{mt},\\ \frac{\mathrm{d}y}{\mathrm{d}t} = \frac{\mathrm{d}x}{\mathrm{d}t} = me^{mt}\\

\implies me^{mt} + e^{mt} - e^{mt}= e^{2t}\\


me^{mt} = e^{2t}\\


me^{m} = e^2 \implies me^m = e^2\\
m\,\textsf{can only be solved for using}\\\textsf{numerical methods, such}\\\textsf{as the Lambert}W\,\textsf{function.}\\

m = W_n(e^2), \hspace{0.1cm} \textsf{where}\,W_n(z) \\\textsf{is the Lambert}W \,\textsf{function}.\\


m = W_n(e^2) \hspace{0.1cm}\textsf{is the solution}\\\textsf{to the equation.}\,\textsf{where}\, n \in \mathbb{Z}\\

\therefore y = x = e^{W_n(e^2)} \hspace{0.1cm}\textsf{is a solution to the first}\\\textsf{order linear ODE.}\, \forall n \in \mathbb{Z}. \\



\textsf{Case 2.}\\

\textsf{If} \, y(t) \,\textsf{is a multiple of}\, x(t) \\


x' + x - y = e^{2t}\\

\displaystyle\frac{\mathrm{d}x}{\mathrm{d}t} + x - y = e^{2t}\\

\textsf{If}\hspace{0.1cm} x = e^{mt}, y = me^{mt} \frac{\mathrm{d}x}{\mathrm{d}t} = me^{mt}\\

\implies me^{mt} + e^{mt} - me^{mt}= e^{2t}\\

e^{mt} = e^{2t} \implies m = 2\\

m = 2\hspace{0.1cm}\textsf{is the only solution}\\\textsf{to the equation.}\\

\therefore x= e^{2t}, y = 2e^{2t} \hspace{0.1cm}\textsf{is a solution to the first}\\\textsf{order linear ODE.}\\

\textsf{We now have two solutions for}\, x \, \& \, y."